OpenStudy (anonymous):

Two Trains A and B starts simultaneously towards each other. One train speed is 5 km/pH more than the other one. After Travelling 4 hrs they are 30 km apart.Find their speed.

OpenStudy (anonymous):



OpenStudy (anonymous):

V1 = x V2 = -(x+5) V12 = x - -(x+5) = 2x +5 S = d/t S*T = d... Hope nw u can solve